# Learning the Hamiltonian Matrix of Large Atomic Systems

### v1

@@ 4.4 Eigenvalue spectrum of predicted a-HfO2
Next, we assess whether the prediction accuracy of the trained network is sufficient for practical application. For this, we use HfO2, as it has the highest error in our experiments in Table 4 and thus presents an upper bound on the expected accuracy. We assemble the full Hamiltonian of the HfO2 test structure using the network outputs (${\bm{H}}^{pred}$), trained with 18 partitions of length $t_{slice}$ = $\sim$3 Å. We then compute the eigenvalue spectrum of ${\bm{H}}^{pred}$ and its reference ${\bm{H}}^{GT}$, as well as the error distribution between them (Fig. 5(a)). The predicted ${\bm{H}}^{pred}$ is able to reproduce all eigenvalues of ${\bm{H}}^{GT}$ within 0.550% relative L1 error. The error decreases to 0.487% when eigenvalues of unoccupied states (those situated above 0.306 $E_{h}$) are excluded. The remaining error is carried mostly by the largest eigenvalues and distributed around the edges of energy gaps (see Fig. 8 in Appendix F), which correspond to regimes of stronger inter-atomic orbital coupling (Atkins & De Paula, 2009).
@@ 5 Conclusion
We developed an equivariant GNN tailored to learn the electronic properties of large-scale, disordered materials, and introduced an augmented partitioning approach to break down and train the graphs encountered when dealing with realistic materials, without sacrificing accuracy. More generally, we proposed a method to tackle the training of systems represented by large, highly connected, and near-sighted graphs where a strictly local environment is sufficient. Our approach can be straightforwardly applied to other complex materials, or adapted to learn their other rotationally-equivariant attributes, such as vibrational properties, e.g., phonon dispersions (Fang et al., 2024). The resulting network captures relevant properties of large structures in sufficient detail to achieve few-$meV$ accuracy and reproduce the energy eigenvalues to under 0.6% error. Further data generation, network optimization, and enabling increased expressiveness will be the next steps.
@@ Appendix B Augmenting graph partitions with virtual nodes
The augmentation approach can be combined with any method of partitioning the input graph. An ideal partitioning scheme would result in the maximal ratio of labeled nodes/edges within each sub-graph, compared to virtual nodes/edges. As we consider structures with fully periodic boundaries in all three dimensions, a simple heuristic to leverage this periodicity is to partition along only one dimension ($x$), thus maintaining all the labeled edges across the $y-$ and $z-$ cell boundaries of each periodic image. This motivates the division by slices described in the main text, which can be very effective when assuming a constant atomic density. Finding more optimal partitions that nevertheless leverage periodicity is a subject of future work.
@@ Appendix C Hyperparameters
We use the hyperparameters shown in Table 6. The ReduceLRonPlaeau scheduler decreases the learning rate by the decay factor when no further decrease in validation loss within the decay patience $t_{patience}$ is detected. The threshold refers to the sensitivity of the scheduler to changes in validation loss. Once the minimum learning rate is reached, the training stops.
@@ Appendix D Dataset generation
Atomic structures corresponding to materials in the amorphous phase are not straightforward to generate since they must accurately capture the structural motifs underlying this phase and a realistic range of atomic coordination. To accurately reproduce long-range structural disorder, the structures used must also be large enough to avoid the creation of wavefunctions that repeat over periodic boundaries. Existing methods to do so include melt-quench (Urquiza et al., 2021), seed-and-coordinate (Youn et al., 2014), or ‘decorate and relax’ (Tafen & Drabold, 2003) approaches. In this work, we use melt-quench processes with molecular dynamics (MD) to evolve each of the three materials considered from their crystalline phases, following a similar procedure as the ones described in (Kaniselvan et al., 2023) and (Urquiza et al., 2021). We then perform a structural relaxation with CP2K code (Kühne et al., 2020) to correct for any discrepancies between the relaxed bond lengths attained with the force field used for MD and those obtained with DFT. Due to the large cell sizes of the a-HfO2 structures, all necessary information is contained within the $\Gamma$ point (where the wavevectors kx = ky = kz = 0). The energies at this location can be computed by directly diagonalizing ${\bm{H}}$. Further details specific to each material are provided in the sections below.
@@
We generate 3 independent structures of a-HfO2 using the QuantumATK toolkit (Søren Smidstrup et al., 2020). As the first step, we run an MD NVT simulation at 3000K for 50 ps with a step size of 1 fs. We use the MTP-HfO2-2022 potential provided by the software. Next, we run an NPT simulation for 300 ps (and the same 1 fs step size), with an initial reservoir temperature of 3000K and a final temperature of 300K, for a cooling rate of 9K/ps. Finally, we anneal the structure at 300K for 50 ps. Due to the computational cost of using a more complete Double-$\zeta$ Valence Polarized (DZVP) basis set (VandeVondele & Hutter, 2007), we use a simpler Single-$\zeta$ Valence (SZV) basis (VandeVondele & Hutter, 2007), which uses 4 basis functions per Oxygen atom and 10 basis functions per Hafnium atom. The plane-wave cutoff is set to 500 Ry, while a cutoff of 60 Ry is used for mapping the Gaussian-type orbitals onto the grid. We use the PBE functional for the exchange-correlation energy (Perdew et al., 1997). To accurately capture the band gap of a-HfO2, we apply the Hubbard correction (Anisimov et al., 1991) of U = 7 eV to the 3d orbital of Ti and the Hubbard correction of U = 10 eV to the 2p orbital of O.
@@ D.2 Substoichiometric HfOx
Two amorphous GST-124 ($Ge(SbTe_{2})_{2}$) structures containing 1008 atoms have been used for training and validation. The first structure is extracted from a crystallization trajectory provided by (Yuxing Zhou, 2023) (Supplementary Material). It is contained in a 25 Å $\times$ 30 Å $\times$ 40 Å orthorhombic bounding box. The second structure is the result of passing a fully crystalline GST-124 structure (the unit cell of which was retrieved from the Materials Project database (Jain et al., 2013) and duplicated to fill a 25 Å $\times$ 30 Å $\times$ 40 Å monoclinic bounding box) through a standard melt-quench procedure (Randomization at 3,000 K (20 ps), cooling to the melting point of 600K at a rate of $10^{14}K.s^{–1}$, holding for 30 ps, quenching to 300 K at $2.5\times 10^{13}K.s^{–1}$ then holding again for another 50 ps). Both structures are relaxed via MD simulations in LAMMPS (Thompson et al., 2022) equipped with the QUIP library for Gaussian Approximation Potential (GAP) (https://github.com/libAtoms/QUIP). The corresponding Hamiltonian terms are obtained using CP2K, where we run the calculations with the DZVP basis, the plane-wave cutoff of 300 Ry, the Gaussian-type orbitals mapping cutoff of 50 Ry, and the PBE functional.
@@
To generate the PtGe structures, germanium structures are taken from the Materials Project database (Jain et al., 2013). This is followed by an NVT melt-quench process using LAMMPS and Stillinger-Weber parameters (Stillinger & Weber, 1985; Nordlund et al., 1998; Wang & Stroud, 1988). The structures are heated to a melting temperature of 5000K at a rate of 0.47 $10^{12}K.s^{–1}$, kept at the melting temperature for 20000 ps (structure 1) or 22000 ps (structure 2), quenched at a rate of 4.7 $10^{12}K.s^{–1}$, and finally annealed at 300K for 100 ps. 1/3 of the Ge atoms are replaced by Pt atoms. The cell of the alloy is then stretched to match the cell of a PtGe2 structure (taken from the Materials Project and optimized using CP2K). Fixed-volume geometry relaxation is then performed on the PtGe alloy. For the structural optimization, as well as for the ${\bm{H}}$ and ${\bm{S}}$ generation, SZV basis set and PBE exchange-correlation functionals are used. We apply a plane-wave cutoff of 1000 Ry and a cutoff for Gaussian-type orbitals mapping of 70 Ry.
@@ Appendix E Atomic bonding environments in the amorphous phase
We use the example of a-HfO2 to investigate the structural differences between samples generated from different starting melt-quench processes. In Fig. 6 we plot the O-coordination of each Hf atom and the radial distribution function $g(r)$ (where r is the inter-atomic distance) for each of the three structures. The distribution in the coordination and dispersion of the peaks in $g(r)$ indicates the amorphous nature of the three structures. Variations between them appear as perturbations in these two quantities. We additionally plot the spatially resolved O-coordination of Hf atoms along the longest, x coordinate for the three structures, as well as the distributions of outliers (Hf atoms with very low and very high O-coordination), demonstrating a significant degree of dissimilarity among the structures.
@@ Appendix F Near-sightedness of the Hamiltonian matrix
Finally, in Fig. 10 we look at the effect of using multiple Message Passing (MP) layers on the training process of HfO2, compared with that of two small molecules (H2O and C3H4O2) taken from the MD17 dataset (Christensen & lilienfeld, 2020). While the use of an additional MP layer has a significant influence on the training of H2O, this effect is reduced for the larger graphs where each atomic environment can contain richer information.
Figure 10: Validation loss as a function of epochs for three structures of varying size (number of atoms). In each case, we compare 1 vs. 2 Message Passing layers. The H2O and C3H4O2 (malondialdehyde) molecules are taken from the MD17 dataset, using train/validate splits of 500/500 molecules. For HfO2 we use an rcut of 8 Å  as in the main text. We do not use augmented partitioning here. Training is performed using a scheduler, and is stopped at 50,000 epochs (H2O and C3H4O2) or 20,000 epochs (HfO2). Data is shown for every 500 epochs.
@@ Appendix G Cutoff radius and connectivity
We now explore the minimal graph connectivity that can be used by the network to accurately learn relevant features of HfO2. To do this we use the slice partition approach introduced in Section 3.3, using a single slice of length $t_{slice}$ = 3 Å  to train the network. Results are reported in (Table. 7). Reducing the value of $r_{cut}$ below 8 Å  noticeably increases the error ($\mathbf{\epsilon_{n/e}}$), thus demonstrating the sensitivity of ${\bm{H}}$ to the exact connectivity of the graph. Going from $r_{cut}$ = 8 Å to 10 Å , the node error begins to plateau, but the node degree (which is proportional to the memory consumption of the network) grows by 1.7$\times$. This is consistent with the negligible changes to the eigenvalue spectra with $r_{cut}$ of 8 Å  (Fig. 8).
$r_{cut}$ [Å] $deg(n)$ $deg(n)^{\prime}$ Epochs $\mathbf{\epsilon_{n}}[mE_{h}]$ $\mathbf{\epsilon_{e}}[mE_{h}]$
4 20.99 10.81 13,816 4.14 9.60
6 74.23 27.78 14,071 3.79 0.40
8 177.09 51.17 18,245 3.76 0.22
10 346.03 81.25 22,463 3.76 0.15
@@
Cutoff [Å] $\mathbf{\epsilon_{n}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{e}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{tot}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{tot}}\left[meV\right]$
6 0.87 0.15 0.17 4.60
8 0.87 0.09 0.10 2.73
16 0.87 0.05 0.05 1.43

### v2

@@ 4.4 Eigenvalue spectrum of predicted a-HfO2
Next, we assess whether the prediction accuracy of the trained network is sufficient for practical application. For this, we use HfO2, as it has the highest error in our experiments in Table 4 and thus presents an upper bound on the expected accuracy. We assemble the full Hamiltonian of the HfO2 test structure using the network outputs (${\bm{H}}^{pred}$), trained with 18 partitions of length $t_{slice}$ = $\sim$3 Å. We then compute the eigenvalue spectrum of ${\bm{H}}^{pred}$ and its reference ${\bm{H}}^{GT}$, as well as the error distribution between them (Fig. 5(a)). The predicted ${\bm{H}}^{pred}$ is able to reproduce all eigenvalues of ${\bm{H}}^{GT}$ within 0.530% relative L1 error. The error decreases to 0.446% when eigenvalues of unoccupied states (those situated above 0.306 $E_{h}$) are excluded. The remaining error is carried mostly by the largest eigenvalues and distributed around the edges of energy gaps (see Fig. 9 in Appendix F), which correspond to regimes of stronger inter-atomic orbital coupling (Atkins & De Paula, 2009).
@@ 4.5 Computational cost
5 Applications in Device Transport Modeling
Investigating materials at the device scale is one of the next challenges to tackle for ML-based materials modeling (Miret et al., 2025). Here we demonstrate that Hamiltonian matrices predicted by our approach can serve as inputs to ab-initio quantum transport simulations, which aim to compute the electrical current flowing through materials under an applied voltage bias. This is done through the solution of the Schrödinger equation with open boundary conditions, where the Hamiltonian matrix contains the electronic properties of the treated device. From the produced energy-resolved transmission function, the electrical current can be derived via the Landauer-Büttiker formula (Brandbyge et al., 2002).
Figure 6: TiN/a-HfO2/Ti/TiN structures with different oxygen vacancy distributions (top), which lead to distinct electrical transmission curves (bottom). The positions of vacant oxygen atoms in the former are indicated with red points. The grey boxes in the transmission plots indicate the difference in the Fermi levels of the two contacts, created through an applied bias. Current flow is determined primarily by the transmission through this energy window.
As a test, we designed a typical metal-insulator-metal resistive switching device by embedding the a-HfOx structure in Fig. 4 with two metal contacts and creating a TiN/a-HfO2/Ti/TiN stack (Kaniselvan et al., 2023). We then computed its electrical current for a-HfOx layers with different vacancy configurations. The electrical transparency of a-HfOx is a function of its exact atomic structure as well as its local stoichiometric variations. Initially, the stoichiometric a-HfO2 blocks current flow due to its finite band gap. The gradual introduction of oxygen vacancies changes this stoichiometry and introduces energy states within the band gap which can carry current (Devices 1-2). This current is further enhanced when defects form a filament bridging the metallic electrodes.
We modeled this process using either a (1) DFT- or (2) ML-generated Hamiltonian. The resulting transmission functions and electrical current values are presented for two TiN/a-HfOx/Ti/TiN structures in Fig. 6 and Table 6, respectively. Training for the latter occurred on slices of amorphous HfOx structures with a uniform distribution of oxygen vacancies. In the test structures, these oxygen vacancies have formed partial or complete conductive filaments. We thus test whether the learned electronic structures of isolated vacancies can be extended to the case of tight interactions between them, which are responsible for the current.
While the quantitative agreement between the transmission functions and electrical currents obtained with a DFT and ML Hamiltonian is not perfect, it is sufficient to (1) track the difference in current across five orders of magnitude, (2) match physically relevant features such as the band edges of the a-HfOx layer or localized regions with higher transmission, and (3) capture the qualitative difference in electrical transparency caused by different spatial distributions of oxygen vacancies. Device simulations with a ML-based Hamiltonian can therefore be used to understand the connection between atomic structure and electrical current flow in TiN/a-HfOx/Ti/TiN stacks. Under this level of accuracy, the model can now also generate Hamiltonian matrices for systems with sizes which are otherwise computationally infeasible with DFT.
Device $\mathbf{\epsilon_{n}}[mE_{h}]$ $\mathbf{\epsilon_{e}}[mE_{h}]$ $I_{ref}$ [A] $I_{pred}$ [A]
0 1.71 0.16 8.00 ×$\times 10^{-9}$ 5.54×$\times 10^{-9}$
1 1.66 0.16 1.48×$\times 10^{-5}$ 1.06×$\times 10^{-5}$
2 1.59 0.15 6.99×$\times 10^{-6}$ 4.89×$\times 10^{-6}$
Table 6: Summary of prediction results and computed currents for devices with vacancy configurations forming different filament shapes. Devices 1-2 are as in Fig. 6.
@@ 6 Conclusion
We developed an equivariant GNN tailored to learn the electronic properties of large-scale, disordered materials, and introduced an augmented partitioning approach to break down and train the graphs encountered when dealing with realistic materials. More generally, we proposed a method to tackle the training of systems represented by large, highly connected, and near-sighted graphs where a strictly local environment is sufficient. Our approach can be straightforwardly applied to other complex materials, or adapted to learn their other rotationally-equivariant attributes, such as vibrational properties, e.g., phonon dispersions (Fang et al., 2024). The resulting network captures relevant properties of large structures in sufficient detail to achieve few-$meV$ accuracy and reproduce the energy eigenvalues to under 0.6% error. Further data generation, network optimization, and enabling increased expressiveness will be the next steps.
Acknowledgements
We acknowledge funding from the ALMOND project (SNSF Sinergia grant no. 198612) and from the Swiss State Secretariat for Education, Research, and Innovation (SERI) through the SwissChips research project. This work was also supported by the Swiss National Supercomputing Center (CSCS) under project lp16. We would also like to thank Paul Uriarte Vicandi and Luiz Felipe Aguinsky for providing the HfO2 and PtGe structures, respectively.
Impact Statement
The goal of this work is to advance research in computational materials science through machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Code and Data Availability
The open source code is available at: https://github.com/chexia8/large_atomic_structures.git.
The datasets used in this work are available at:
https://huggingface.co/datasets/chexia8/Amorphous-Hamiltonians/.
@@ Appendix B Augmenting graph partitions with virtual nodes
In Algorithm 1, we detail the procedure to partition the full graph ${\mathcal{G}}$, described by the set of vertices $\mathcal{V}$ and edges $\mathcal{E}$, into a set of slices $\{{\mathcal{G}}_{1}\;\ldots\;{\mathcal{G}}_{N}\}$ which are augmented by virtual nodes and edges. The number of labeled nodes ($n_{i}^{n}$) and the number of labeled edges ($n_{i}^{e}$) are collected and passed to the training functions, which then omit the remainder of the outputs (the virtual nodes and edge outputs) while computing the loss.
1
Input: Graph ${\mathcal{G}}(\mathcal{V},\mathcal{E})$
Input: Slice x-coordinates $[{x}_{1}\;\ldots\;{x}_{N}]$
Output: Set of subgraphs $\{{\mathcal{G}}_{1}\;\ldots\;{\mathcal{G}}_{N}\}$
Output: Numbers labeled nodes $[n^{n}_{1}\;\ldots\;n^{n}_{N}]$
Output: Numbers labeled edges $[n^{e}_{1}\;\ldots\;n^{e}_{N}]$
2
3for $i\leftarrow 1$ to $N$ do
[…]
5       $n^{n}_{i}=0$;
6       for $v\in\mathcal{V}$ do
7             if $v.x\in[x_{i},x_{i+1})$ then
[…]
9                   $n^{n}_{i}\mathrel{+}=1$;
10
11             end if
12
13       end for
[…]
15       $n^{e}_{i}=0$;
16       for $v_{1}\in\mathcal{V}_{i}$ do
17             for $v_{2}\in\mathcal{V}_{i}$ do
18                   if $v_{1}\rightarrow v_{2}\in\mathcal{E}$ then
[…]
20                         $n^{e}_{i}\mathrel{+}=1$;
21
22                   end if
23
24             end for
25
26       end for
27      for $v_{1}\in\textnormal{{$\mathcal{V}$}}_{i}$ do
28             for $v_{2}\in\mathcal{V}\setminus\textnormal{{$\mathcal{V}$}}_{i}$ do
29                   if $v_{2}\rightarrow v_{1}\in\mathcal{E}$ then
[…]
32
33                   end if
34
35             end for
36
37       end for
[…]
39
40 end for
Algorithm 1 Augmented partitioning approach
The augmentation approach can be combined with any method of partitioning the input graph. An ideal partitioning scheme would result in the maximal ratio of labeled nodes/edges within each sub-graph, compared to virtual nodes/edges. As we consider structures with fully periodic boundaries in all three dimensions, a simple heuristic to leverage this periodicity is to partition along only one dimension ($x$), thus maintaining all the labeled edges across the $y-$ and $z-$ cell boundaries of each periodic image. This motivates the division by slices described in Algorithm 1, which can be very effective when assuming a constant atomic density. Finding more optimal partitions that nevertheless leverage periodicity is a subject of future work.
@@ Appendix C Hyperparameters
We use the hyperparameters shown in Table 7 to train on all datasets. The ReduceLRonPlaeau scheduler decreases the learning rate by the decay factor when it does not detect a further decrease in validation loss within the decay patience $t_{patience}$. The threshold refers to the sensitivity of the scheduler to changes in validation loss. Once the minimum learning rate is reached, the training stops.
@@ Appendix D Dataset generation
Atomic structures corresponding to materials in the amorphous phase are not straightforward to generate since they must accurately capture the structural motifs underlying this phase and a realistic range of atomic coordination. To accurately reproduce long-range structural disorder, the structures used must also be large enough to avoid the creation of wavefunctions that repeat over periodic boundaries. Existing methods to do so include melt-quench (Urquiza et al., 2021), seed-and-coordinate (Youn et al., 2014), or ‘decorate and relax’ (Tafen & Drabold, 2003) approaches.
In this work, we use melt-quench processes with molecular dynamics (MD) to evolve each of the three materials considered from their crystalline phases, following a similar procedure as the ones described in (Kaniselvan et al., 2023) and (Urquiza et al., 2021). We then perform a structural relaxation with CP2K code (Kühne et al., 2020) to correct for any discrepancies between the relaxed bond lengths attained with the force field used for MD and those obtained with DFT. Due to the large cell sizes of the a-HfO2 structures, all necessary information is contained within the $\Gamma$ point (where the wavevectors kx = ky = kz = 0). The energies at this location can be computed by directly diagonalizing ${\bm{H}}$. The datasets are publicly available at https://huggingface.co/datasets/chexia8/Amorphous-Hamiltonians.
Further details specific to each material are provided in the sections below.
@@
We generate 3 independent structures of a-HfO2 using the QuantumATK toolkit (Søren Smidstrup et al., 2020). As a first step, we run an MD NVT simulation at 3000K for 50 ps with a step size of 1 fs. We use the MTP-HfO2-2022 potential provided by the software. Next, we run an NPT simulation for 300 ps (and the same 1 fs step size), with an initial reservoir temperature of 3000K and a final temperature of 300K, for a cooling rate of 9K/ps. Finally, we anneal the structure at 300K for 50 ps.
Due to the computational cost of using a more complete Double-$\zeta$ Valence Polarized (DZVP) basis set (VandeVondele & Hutter, 2007), we use a simpler Single-$\zeta$ Valence (SZV) basis (VandeVondele & Hutter, 2007), which uses 4 basis functions per Oxygen atom and 10 basis functions per Hafnium atom. The plane-wave cutoff is set to 500 Ry, while a cutoff of 60 Ry is used for mapping the Gaussian-type orbitals onto the grid. We use the PBE functional for the exchange-correlation energy (Perdew et al., 1997). To accurately capture the band gap of a-HfO2, we apply the Hubbard correction (Anisimov et al., 1991) of U = 7 eV to the 3d orbital of Ti and the Hubbard correction of U = 10 eV to the 2p orbital of O.
@@ D.2 Substoichiometric HfOx
We use two amorphous GST-124 ($Ge(SbTe_{2})_{2}$) structures containing 1008 atoms for training and validation. The first structure is extracted from the MD simulation of GST-124 crystallization, provided by (Yuxing Zhou, 2023) (Supplementary Material). The initial geometry for the second structure is a perfectly crystalline GST-124 structure, taken from Materials Project database (Jain et al., 2013). The amorphous structure is then generated through a standard melt-quench procedure, consisting of atomic position randomization at 3000K for 20 ps, cooling to the melting point of 600K at a rate of $10^{14}K.s^{–1}$, holding for 30 ps, quenching to 300K at a rate of $2.5\times 10^{13}K.s^{–1}$, and finally, annealing at 300K for 50 ps. Both amorphous structures are obtained via MD simulations in LAMMPS (Thompson et al., 2022), equipped with the QUIP library for Gaussian Approximation Potential (GAP) (Csányi et al., 2007). The corresponding Hamiltonian matrices are obtained using CP2K, where we run calculations with the DZVP basis, the plane-wave cutoff of 300 Ry, the Gaussian-type orbitals mapping cutoff of 50 Ry, and the PBE functional.
@@
To generate the PtGe structures, germanium structures are taken from the Materials Project database (Jain et al., 2013). This is followed by an NVT melt-quench process using LAMMPS and Stillinger-Weber parameters (Stillinger & Weber, 1985; Nordlund et al., 1998; Wang & Stroud, 1988). The structures are heated to a melting temperature of 5000K at a rate of 0.47 $10^{12}K.s^{–1}$, kept at the melting temperature for 20000 ps (structure 1) or 22000 ps (structure 2), quenched at a rate of 4.7 $10^{12}K.s^{–1}$, and finally annealed at 300K for 100 ps. Using the Atomic Simulation Environment (ASE) tool (Hjorth Larsen et al., 2017), 1/3 of the Ge atoms are replaced by Pt atoms. The cell of the alloy is then stretched to match the cell of a PtGe2 structure (taken from the Materials Project and optimized using CP2K). Fixed-volume geometry relaxation is then performed on the PtGe alloy. For the structural optimization, as well as for the ${\bm{H}}$ and ${\bm{S}}$ generation, SZV basis set and PBE exchange-correlation functionals are used. We apply a plane-wave cutoff of 1000 Ry and a cutoff for Gaussian-type orbitals mapping of 70 Ry.
@@ Appendix E Atomic bonding environments in the amorphous phase
We use the example of a-HfO2 to investigate the structural differences between samples generated from different starting melt-quench processes. In Fig. 7 we plot the O-coordination of each Hf atom and the radial distribution function $g(r)$ (where r is the inter-atomic distance) for each of the three structures. The distribution in the coordination and dispersion of the peaks in $g(r)$ indicates the amorphous nature of the three structures. Variations between them appear as perturbations in these two quantities. To gain more insights into how different the structures are, we additionally plot the spatially resolved O-coordination of Hf atoms along the longest, x coordinate for the three structures, as well as the distributions of outliers (Hf atoms with very low and very high O-coordination) in three-dimensional space. These outliers are situated at different locations in different structures, demonstrating a significant degree of dissimilarity among the structures.
@@ Appendix G Cutoff radius and connectivity
We now explore the minimal graph connectivity that can be used by the network to accurately learn relevant features of HfO2. To do this we use the slice partition approach introduced in Section 3.3, using 18 slices of length $t_{slice}$ = 3 Å  to train the network. Results are reported in (Table. 8). Reducing the value of $r_{cut}$ below 8 Å  noticeably increases the error ($\mathbf{\epsilon_{n/e}}$), thus demonstrating the sensitivity of ${\bm{H}}$ to the exact connectivity of the graph. Going from $r_{cut}$ = 8 Å to 12 Å , the edge error improves, while the node error begins to plateau, but the node degree (which is proportional to the memory consumption of the network) grows by 1.7$\times$. This is consistent with the negligible changes to the eigenvalue spectra with $r_{cut}$ of 8 Å  (Fig. 9).
$r_{cut}$ [Å] $deg(n)$ $deg(n)^{\prime}$ $\mathbf{\epsilon_{n}}[mE_{h}]$ $\mathbf{\epsilon_{e}}[mE_{h}]$ $\mathbf{\epsilon_{tot}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{tot}}\left[meV\right]$
4 21.02 10.60 2.84 0.52 0.64 17.50
6 72.41 32.01 2.27 0.26 0.29 7.98
8 177.61 49.89 2.16 0.16 0.18 4.84
10 347.16 78.85 2.18 0.12 0.12 3.38
12 590.25 138.59 2.13 0.09 0.10 2.58
@@
Cutoff [Å] $deg(n)$ $deg(n)^{\prime}$ $\mathbf{\epsilon_{n}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{e}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{tot}}\left[mE_{h}\right]$ $\mathbf{\epsilon_{tot}}\left[meV\right]$
4 14.28 10.00 1.08 0.26 0.33 8.96
6 50.17 27.37 0.80 0.14 0.16 4.28
8 118.71 51.57 0.77 0.08 0.09 2.49
10 234.28 84.07 0.80 0.08 0.08 2.17